# AutoAct: Automatic Agent Learning from Scratch via Self-Planning

### 10 Jan 2024 v1

@@ 7 Conclusion
In this paper, we propose AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models, while alleviating the pressure on individual agents by explicitly dividing the workload, thereby enhancing the collective performance of multi-agents. Experimental results demonstrate that AutoAct performs superior on challenging question-answering benchmarks compared to various strong baselines, and AutoAct with Llama-2-13b model can even obtain comparable performance of GPT-3.5-Turbo agent. Extensive analysis indicates the effectiveness of our appropriate division-of-labor strategy, with the trajectory quality generated by AutoAct significantly outperforming that of other methods from multiple aspects.
@@
Benchmarks.
For experimental convenience, we only evaluate our method on complex question-answering tasks. However, there are many other interactive scenarios beyond this, such as web Yao et al. (2022); Zhou et al. (2023b), household Puig et al. (2018); Shridhar et al. (2021), robotics Ichter et al. (2022), and so on. We plan to conduct further research on applying AutoAct to a wider range of benchmarks in the future.
Boosting Knowledge through Self-Instruct.
@@ Appendix B Task Information
Question: Who directed the Academy Award-winning film Giorgio Cantarini appeared in in 2000?
Answer: Ridley Scott
@@
Question: Which company, that had Greg McAdoo on their board, was acquired by Google?
Answer: Bump Technologies
Question: What is the ethnicity of Ivanka Trump’s mother?
Answer: Czech-American
@@
: This is a multimodal question-answering task that necessitates a language model to utilize tools for transforming image information into textual data. Simultaneously, this task incorporates substantial background knowledge, requiring the language model to acquire external information to enhance its comprehension of the task.
@@
Question: Identify the question that Tom and Justin’s experiment can best answer. Context: The passage below describes an experiment. Read the passage and then follow the instructions below. Tom placed a ping pong ball in a catapult, pulled the catapult’s arm back to a 45 angle, and launched the ball. Then, Tom launched another ping pong ball, this time pulling the catapult’s arm back to a 30 angle. With each launch, his friend Justin measured the distance between the catapult and the place where the ball hit the ground. Tom and Justin repeated the launches with ping pong balls in four more identical catapults. They compared the distances the balls traveled when launched from a 45 angle to the distances the balls traveled when launched from a 30 angle. Figure: a catapult for launching ping pong balls.
@@
Question: What is the mass of a pair of goggles? Context: Select the better estimate.
Options: (A) 60 grams (B) 60 kilograms
Caption: no img
Answer: A. 60 grams
Question: Which figure of speech is used in this text? Sing, O goddess, the anger of Achilles son of Peleus, that brought countless ills upon the Achaeans.Homer, The Ilia
Options: (A) chiasmus (B) apostrophe
Caption: no img
Answer: B. apostrophe
Question: Which country is highlighted?
Options: (A) Dominica (B) Haiti (C) Cuba (D) The Bahamas
Caption: A person flying through the air with a kite.
Answer: B. Haiti
@@ Appendix C Tool Library
BingSearch[query], which searches the exact detailed query on the Internet and returns the relevant information to the query. Be specific and precise with your query to increase the chances of getting relevant results. For example,;Bingsearch[popular dog breeds in the United States]

### 17 Jan 2024 v2

@@ 7 Conclusion
In this paper, we propose AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models, while alleviating the pressure on individual agents by explicitly dividing the workload, thereby enhancing the collective performance of multi-agents. Experimental results demonstrate that AutoAct performs superior on challenging question-answering benchmarks compared to various strong baselines, and AutoAct with Llama-2-13b model can even obtain comparable performance of zero-shot GPT-3.5-Turbo agent. Extensive analysis indicates the effectiveness of our appropriate division-of-labor strategy, with the trajectory quality generated by AutoAct significantly outperforming that of other methods from multiple aspects.
@@
Tasks and Backbones.
For experimental convenience, we only evaluate our method on complex question-answering tasks with the Llama-2-chat model series. However, there are many other interactive scenarios and backbone models beyond these. Other complex tasks include web Yao et al. (2022); Zhou et al. (2023b), household Puig et al. (2018); Shridhar et al. (2021), robotics Ichter et al. (2022), etc., and more backbone models include Vicuna Zheng et al. (2023), ChatGLM Du et al. (2022), Mistrial Jiang et al. (2023), etc. We plan to conduct further research on applying AutoAct to a wider range of tasks and backbones in the future.
Boosting Knowledge via Self-Instruct.
@@ Appendix B Task Information
: This is a multimodal question-answering task that necessitates a model to utilize tools for transforming image information into textual data. Simultaneously, this task incorporates substantial background knowledge, requiring the language model to acquire external information to enhance its comprehension of the task.
@@
Question: Identify the question that Tom and Justin’s experiment can best answer.
Context: The passage below describes an experiment. Read the passage and then follow the instructions below. Tom placed a ping pong ball in a catapult, pulled the catapult’s arm back to a 45 angle, and launched the ball. Then, Tom launched another ping pong ball, this time pulling the catapult’s arm back to a 30 angle. With each launch, his friend Justin measured the distance between the catapult and the place where the ball hit the ground. Tom and Justin repeated the launches with ping pong balls in four more identical catapults. They compared the distances the balls traveled when launched from a 45 angle to the distances the balls traveled when launched from a 30 angle. Figure: a catapult for launching ping pong balls.
@@ Appendix C Tool Library
BingSearch[query], which searches the exact detailed query on the Internet and returns the relevant information to the query. Be specific and precise with your query to increase the chances of getting relevant results. For example, Bingsearch[popular dog breeds in the United States]